OpenStudy (anonymous):

There are 16 girls and 12 boys on the cross country team

OpenStudy (escamer):

then?

OpenStudy (theedwardsfamily):

Where's the rest?

OpenStudy (anonymous):

theres just 16 girls and 12 boys thats it

OpenStudy (escamer):

there is no problem then

OpenStudy (anonymous):

guess not

OpenStudy (anonymous):

There are 16 girls and 12 boys on the cross country team. What fraction of the team is made up of boys?

OpenStudy (escamer):

12/16

OpenStudy (ihelpyoulearn):

3/7! For sure

OpenStudy (escamer):

simplifying it is 3/4 i think

OpenStudy (theedwardsfamily):

Simplest form right?

OpenStudy (anonymous):

yeah

OpenStudy (ihelpyoulearn):

3/7! You add the boys and girls up which is 28 so 12/28=3/7

OpenStudy (anonymous):

First, you have to add 16 and 12 to gain the total amount of kids on the team. 16 + 12 = 28 so... there are 28 kids in total; 16 of which are girls and 12 of which are boys. The answer is easy: 12/28 Why? Because out of 28 kids, 12 are boys. Now you have to simplify your answer. The best number to simplify with would be 4. 12 ÷ 4 = 3 28 ÷ 4 = 7 So 3/7 is the fraction of boys on the cross country team as well as your final answer.

OpenStudy (ihelpyoulearn):

Great response @LYRA_Z!